OpenStudy (anonymous):

A coin collection jar is full of just nickels and dimes. If there are 355 coins valued at $24.45, how many nickels are in the jar?

OpenStudy (whpalmer4):

two equations: # of nickels + # of dimes = 355 # of nickels * 0.05 + # of dimes * 0.10 = 24.45 solve the system and you'll have your answer

OpenStudy (anonymous):

the important thing is that you have 24.45 dollar. that's all matter :D

OpenStudy (whpalmer4):

do you know how to solve such a system of equations?

OpenStudy (anonymous):

Yes, give me a minute, I was trying to figure it out and didn't realize I got a reply on here.

OpenStudy (anonymous):

I'm stuck..

OpenStudy (johnweldon1993):

What are you stuck on?

OpenStudy (whpalmer4):

what do you have for the equations?

OpenStudy (anonymous):

I don't understand the equation that was given up there^

OpenStudy (anonymous):

isn't there supposed to be variables?

OpenStudy (johnweldon1993):

Lets let nickels be "n" and Dimes be "d" We know there are some amount of both that add to 355 coins... Now we also know that each nickel is .05 and each dime is .10...this adds up to 24.45 so n + d = 355 .05n + .10d = 24.45 how's that look?

OpenStudy (whpalmer4):

well, I gave them in words: "# of dimes" could be d, # of nickels could be n

OpenStudy (anonymous):

okay yeah i wrote down 0.05n+0.10d=24.25?

OpenStudy (anonymous):

now i solve for n?

OpenStudy (whpalmer4):

not quite: 24.45!

OpenStudy (johnweldon1993):

well 24.45 but yeah

OpenStudy (anonymous):

oh, typo lol :)

OpenStudy (whpalmer4):

typo which will destroy any chance at a correct answer...

OpenStudy (anonymous):

lol! okay I got: n=24.45-0.10d/0.05?

OpenStudy (whpalmer4):

here's how I would do it: n + d = 355 0.05n + 0.10d = 24.45 multiply the second equation by 100 to clear out the pesky decimals n+d = 355 5n + 10d = 2445 (alternatively, think of the problem in cents instead of dollars) now solve the first equation for either n or d, and plug that in the second one

OpenStudy (whpalmer4):

giving you an equation in just one variable which is easily solved. then use the first equation to get the value of the other variable

OpenStudy (anonymous):

wait, first equation as in n+d=355?

OpenStudy (anonymous):

n=355-d

OpenStudy (johnweldon1993):

Right...so now substitute that into your second equation for 'n'

OpenStudy (anonymous):

355-n=2445-10d/5 and solve for d?

OpenStudy (johnweldon1993):

5n + 10d = 2445 after plugging in n = 355 - d 5(355 - d) + 10d = 2445 and simplify from here...do you see what I did there? simply replaced 'n' with what 'n' equals (355 - d)

OpenStudy (anonymous):

ohhh, ahh i did it wrong again. hold on

OpenStudy (anonymous):

d=134 n=221?

OpenStudy (johnweldon1993):

That is what I got!

OpenStudy (anonymous):

i plugged the amount of dimes back into 5n+10d=2445 right?

OpenStudy (johnweldon1993):

Yeah recheck your answer....plug both those numbers back in and see if you get 2445

OpenStudy (anonymous):

Yes sir, 5(221)+10(134)=2445

OpenStudy (johnweldon1993):

So then we have the correct answer :)

OpenStudy (anonymous):

Thank you :) I'm horrible with word problems lol

OpenStudy (anonymous):

Thank you to you to @whpalmer4 :)

OpenStudy (whpalmer4):

in general, you need to check your solution in all of the equations of the system: it's possible to come up with "solutions" that work for only some of the equations, not all. for example, 240 dimes and 9 nickels would give you $24.45, but fail the 355 coins equation

OpenStudy (anonymous):

wait thats wrong ;/ it says 225 is nickels

OpenStudy (anonymous):

and its wrong by my mistake, dang it! it's $24.25 in the book.

OpenStudy (anonymous):

oh lordy, let me see if i can fix it though and get the right answer

OpenStudy (johnweldon1993):

Ahh...well yes if it is 24.25 in the book then yes 225 would be correct...but yeah lol

OpenStudy (anonymous):

okay got it! d=130 and nickels=225

OpenStudy (anonymous):

=2425

OpenStudy (whpalmer4):

I'm going to give you two more ways you could solve this: 1-variable approach: if there are 355 coins, we can say that x is the number of nickels, and 355-x is the number of dimes, right? we could write just one equation: 5x + 10(355-x) = 2425 (I'm doing it in cents to skip the multiply by 100 step) 5x + 3550 - 10x = 2425 3550 - 5x = 2425 3550-2425 = 5x 1125 = 5x 225 = x so 225 nickels, 355-225 = 130 dimes in your head approach: assume all of the coins are dimes: 355 dimes = $35.50. that's too high by $35.50-$24.25 = $11.25. each dime we change to a nickel reduce the amount by $0.05. We need $11.25/$0.05 = 225 nickels instead of dimes.

OpenStudy (whpalmer4):

if you have an easy way of making graphs, graphing the lines n+d=355 and 5n+10d=2425 and finding the point where they intersect also gives you the solution.

OpenStudy (anonymous):

thank you :)

OpenStudy (whpalmer4):

it turns out that most word problems you commonly encounter fall into a couple of different categories, and once you figure out to solve the problems in each category, you're in good shape. there is a good book by Rebecca Wingard-Nelson that you can probably find at the library and become a wizard at doing these :-)